### v1

# Cross-Modal Geometric Hierarchy Fusion: An Implicit-Submap Driven Framework for Resilient 3D Place Recognition

Abstract
LiDAR-based place recognition serves as a crucial enabler for long-term autonomy in robotics and autonomous driving systems. Yet, prevailing methodologies relying on handcrafted feature extraction face dual challenges: (1) Inconsistent point cloud density, induced by ego-motion dynamics and environmental disturbances during repeated traversals, leads to descriptor instability, and (2) Representation fragility stems from reliance on single-level geometric abstractions that lack discriminative power in structurally complex scenarios. To address these limitations, we propose a novel framework that redefines 3D place recognition through density-agnostic geometric reasoning. Specifically, we introduce an implicit 3D representation based on elastic points, which is immune to the interference of original scene point cloud density and achieves the characteristic of uniform distribution. Subsequently, we derive the occupancy grid and normal vector information of the scene from this implicit representation. Finally, with the aid of these two types of information, we obtain descriptors that fuse geometric information from both bird’s-eye view (capturing macro-level spatial layouts) and 3D segment (encoding micro-scale surface geometries) perspectives. We conducted extensive experiments on numerous datasets (KITTI, KITTI-360, MulRan, NCLT) across diverse environments. The experimental results demonstrate that our method achieves state-of-the-art performance. Moreover, our approach strikes an optimal balance between accuracy, runtime, and memory optimization for historical maps, showcasing excellent Resilient and scalability. Our code will be open-sourced in the future.
Index Terms:
LiDAR, place recognition, Implicit Neural Representation, 3D descriptors
I Introduction
PLACE recognition, the capability of identifying previously visited locations in unmanned systems, serves as a cornerstone for achieving long-term autonomy in robotics and self-driving platforms [1]. This capability is crucial for autonomous mobile robots to achieve precise and robust positioning in unknown environments [2]. Place recognition also has a wide range of applications. For instance, it is used in loop closure detection for Simultaneous Localization and Mapping (SLAM) [3, 4, 5], map merging in large-scale scene mapping [6], and autonomous robot navigation [7, 8]. Light Detection and Ranging (LiDAR) can obtain accurate measurement results even under various lighting conditions, such as in darkness or when disturbed by strong light. As a result, LiDAR plays a crucial role in unmanned systems. Meanwhile, with the rapid development of robotic and autonomous driving systems, there is an increasing demand for LiDAR-based place recognition[9, 10, 11, 12].
Despite the rapid development of place recognition methods based on LiDAR point clouds in recent years, these methods still face the following challenges:
  • $\bullet$
    In an urban environment, the LiDAR point clouds collected at the same location often exhibit inconsistencies in density and coverage due to the variability in the trajectories and motion speeds of vehicles or robots. Therefore, the algorithm must possess variability robust rotational and translational invariance, as well as invariance to point cloud density[13, 14].
  • $\bullet$
    Dynamic objects within the scene, such as cars, pedestrians, and cyclists, continuously alter the 3D structure and distribution of the environment. Their presence is unavoidable and challenges the robustness of LiDAR point cloud-based place recognition [15, 16].
Figure 1: Our 3D place recognition pipeline, unlike other methods, first converts the submaps into implicit neural point representations. The colored dots in the graph represent these neural representation points. Subsequently, utilizing the advantages of implicit representations, we fuse information from two modalities, namely BEV features and surface normal information of 3D segments, to generate descriptors for each submap.
Most existing methods for hand-crafted feature-extraction-based 3D place recognition operate in two primary strategies. One approach is to extract features (such as local[17, 18] or global features[19, 20]) directly from the original 3D point cloud. The other approach involves first converting the raw point cloud into a structured representation (for example, using Bird’s Eye View (BEV)[21], polar coordinate projection[22], intensity images[23], or sparse voxel grids[24]) and then extracting features.
However, existing 3D-based hand-crafted feature extraction methods generally suffer from a fundamental limitation: their reliance on relatively simplistic scene representations. These methods typically focus on feature extraction from single-modal scene representations—either directly from raw point cloud data or by converting point clouds into a specific structured representation (such as BEV or polar coordinates) and relying on it. This dependence on single representations inevitably leads to incompleteness of scene information. The reason is that while raw point clouds preserve precise 3D geometric details, they are susceptible to density variations and lack inherent rotational and translational invariance; structured representations (e.g., BEV), although providing compact global perspectives conducive to performing feature extraction, often sacrifice fine-resolution local geometry or vertical-direction information. Consequently, it is difficult to use features extracted from any single representation to comprehensively and robustly characterize complex 3D scenes, thereby constraining the upper limit of performance for subsequent 3D place recognition tasks.
To address these issues, we have deviated from the previous explicit 3D point cloud input and introduced a point-based implicit representation as the input data. We introduce this representation for three reasons:
  • $\bullet$
    The point-based implicit representation is compact, flexible, and continuous. It possesses rotational and translational invariance, as well as spatial density invariance.
  • $\bullet$
    This representation can efficiently provide high-quality, globally consistent surface normals, offering the opportunity to derive more discriminative place recognition descriptors.
  • $\bullet$
    This representation enables more flexible online removal of dynamic objects, leading to more robust place recognition.recognition.
This paper proposes a brand-new place recognition method based on LiDAR point clouds, as illustrated in Fig. 1. This method uses a point-based implicit representation as the input for scene representation and utilizes the Signed Distance Function (SDF) and stability prediction values to filter out online dynamic objects. Subsequently, it integrates normal and BEV information to enhance place recognition capability. Specifically, we construct a normal-based descriptor by leveraging the angular differences between the surface normals of the key 3D segment. Meanwhile, we generate evenly distributed occupancy grid information using the point-based implicit representation and compute BEV descriptors. Finally, we combine these two types of descriptors to form the final descriptor.
In summary, the contributions of this article are as follows:
  • $\bullet$
    We introduce a point-based implicit representation. Leveraging its consistent properties and high-quality surface normals, we construct a novel point cloud-based place recognition scheme.
  • $\bullet$
    We propose a cross-modal geometric hierarchy fusion method that synthesizes structural normal descriptors with rotation-invariant BEV features, enabling resilient 3D place recognition through complementary geometric representations.
  • $\bullet$
    We propose a training-free, handcrafted feature extraction method. Our implicit representation enables substantial storage savings and real-time operation. Extensive experiments demonstrate that our method outperforms state-of-the-art approaches.
II Related Works
In this section, we will provide a comprehensive review of implicit map representation, as well as 3D point cloud place recognition methods based on handcrafted and deep learning.
II-A Implicit Neural Map Representation
In recent decades, maps based on explicit representation have been widely applied in unmanned systems[25, 26]. Common methods for explicitly representing scenes include point clouds[27], triangular meshes[28], voxel grids[29], occupancy grids[30], and so on. Recently, the effectiveness of implicit neural representation has been demonstrated in the modeling of radiation fields[31] and geometric fields[32, 33]. Such a representation has facilitated many application scenarios, such as 3D reconstruction, scene generation, novel view synthesis, and so on. Based on NeRF (Neural Radiance Field)[31], DeepSDF[32], and occupancy networks[33], a large amount of research work has been carried out. However, all these works use only a single multi-layer perceptron (MLP) to represent the entire scene. Although this approach is effective, it may face poor scalability, long computation time, and high memory consumption when dealing with complex scenes. Therefore, recent approaches have adopted the hybrid representation to address the above-mentioned issues. This approach jointly optimizes the explicitly stored local latent features and the shallow multi-layer perceptron[34, 35, 36].
With remarkable progress achieved in efficient training methods for implicit neural representations, it has become possible to scale them up to larger scenarios, and they have been widely applied in mapping and Simultaneous Localization and Mapping (SLAM). In the mapping and SLAM systems based on RGB-D data streams, researchers have mainly proposed two types of methods. One is based on a single MLP[37, 38], and the other is based on grid-based local latent features and shallow MLPs[39, 40, 41] to model the geometry or radiance field of the scene while tracking the camera’s pose.
Some work has also been done for the cases where LiDAR is used as the input data stream. IR-MCL[42] and LocNDF[43] proposed a method to construct an implicit neural distance map through laser scanning and to localize the robot within this map. SHINE-mapping[44] aims to extend the implicit neural representation to large-scale outdoor LiDAR data. It stores local features in a sparse voxel format based on an octree, enabling it to handle large-scale scene data while maintaining efficient storage. In the application of SLAM, LONER[45] integrates incremental neural mapping with an ICP-based LiDAR odometry front-end. This approach combines the modeling capabilities of neural networks with traditional point cloud registration algorithms to enhance the accuracy and efficiency of map construction and localization. NeRF-LOAM[46] proposes an implicit neural LiDAR odometry and mapping system based on an online optimizable octree-based feature grid. This method utilizes the concept of NeRF and improves the quality of map construction and the accuracy of localization by optimizing the feature grid. Compared with the previous systems, PIN-SLAM[47] has achieved greater improvements in real-time performance and map storage efficiency. It represents the scene using elastic neural points with fast indexing based on a hash table, enabling the system to operate at a speed close to real-time and saving storage overhead. Inspired by the above work, our work employs a point-based implicit representation. From this representation, we obtain the scene normal vector and occupancy information, which are the foundation for subsequent place recognition.
II-B Handcrafted Methods
Early studies mainly described point clouds by extracting their geometric local features, including distance[48], point normal[49], angle[50], density[51], and so on. These methods have achieved success in point cloud registration and shape recognition. With the in-depth development of the research, researchers have started to attempt to encode the entire point cloud into a global feature descriptor. Z-Projection[52] projects the point cloud onto a certain direction (such as the Z-axis), extracts angular and height information, and then calculates the similarity through the Sørensen distance or the Wasserstein distance. Fast Histogram[53] uses a fast histogram to statistically analyze angular and height information and evaluates the similarity through the Wasserstein distance. M2DP[54] projects the point cloud onto multiple two-dimensional planes, calculates the point density on each plane, and forms a signature vector. This method avoids the calculation of point normals, improving the computational efficiency, but some information may be lost. Scan Context[55, 56] divides the horizontal space into multiple annular regions and fan-shaped regions, forming a two-dimensional height matrix. By performing nearest neighbor search through the annular key and the similarity score, good results have been achieved. To further enhance the performance, researchers have started to combine other types of features, such as intensity[57], spatial binary patterns[58], and so on. LiDAR Iris[59] converts point cloud data into ”iris images”, generates binary signatures through the LoG-Gabor filter and threshold operations, and then calculates the similarity through the Hamming distance. STD[18] proposes a stable triangle descriptor based on the invariance of the triangle shape. By extracting key points and encoding them into triangle descriptors, it enables efficient position matching and geometric verification. Most of the handcrafted methods are relatively natural. However, they generally encounter some problems when faced with large rotations and translations.
II-C Learning-based Methods
PointNetVLAD[17] is a deep learning-based method that combines two crucial components: PointNet[60] and NetVLAD[61]. PointNet is a network that extracts features from point clouds, while NetVLAD is an aggregator that aggregates local features into a 512-dimensional global descriptor. This method lays the foundation for subsequent approaches. Based on PointNetVLAD, two methods, namely SOE-Net[62] and PCAN[63], introduce attention mechanisms. The attention mechanisms can dynamically emphasize the more important parts in the point cloud, thereby enhancing the robustness of the descriptors. Locus[64] enhances the descriptors by introducing temporal information, which can improve the robustness against viewpoint changes. HiTPR[65] partitions the point cloud into dense voxels and utilizes a transformer network[66] to enhance the correlation of local neighbors and the dependence on global context. Consequently, this method demonstrates outstanding performance in highly complex and cluttered environments. Kim et al.[67] proposed a classifier based on Convolutional Neural Network (CNN) for long-term place recognition. This approach can handle significant scene changes over a long time. Semantic Scan Context[68] addresses the translation problem by introducing semantic information. LPD-Net[69] and vLPD-Net[70] utilize Graph Neural Network (GNN) for feature aggregation, thereby effectively capturing the geometric and shape characteristics of the point cloud. DAH-Net[71] proposes a density-driven adaptive hybrid network based on the density changes of point clouds. Through a dynamic local feature aggregation and a contrastive enhanced linear attention module, it attempts to address the challenges posed by the density variations of point clouds in large-scale scenes. These aforementioned learning-based methods often suffer from relatively long training times and require substantial computational resources. Moreover, they exhibit poor performance when faced with scenarios that are not part of their training samples.
III METHOD
Figure 2: Here is the pipeline of our 3D Place Recognition. A specific number of LiDAR frames are fed into form a sub-map, which is then transformed into an implicit neural point representation. High-quality occupancy grid information, triangular points, and normal vector information are subsequently derived from this representation. BEV and primary 3D segments are then obtained. The log-Gabor filter is employed to generate descriptors for the BEV. At the same time, the angular difference distribution of normal vectors is utilized to obtain geometric information descriptors for 3D segments. Finally, these two types of descriptors are fused to create a descriptor for 3D place recognition.
This section will introduce our proposed method for Lidar-based place recognition. In Section III-A, an overview of our method will be presented. Section III-B will demonstrate the neural implicit representation. Section III-D and Section III-E will describe how we obtain descriptors with normal and BEV image information.
III-A Overview
The main process of our proposed method is illustrated in Fig. 2. To better handle the random noise caused by variations in point cloud density and dynamic objects and obtain more discriminative descriptors, we use elastic implicit 3D points to represent the scene.
Considering that the LiDAR point clouds in different scanning modes need to be processed, the input of our method is several frames of point clouds $\{\mathbf{P_{1}},...,\mathbf{P_{n}}\}$. The value of $n$ is determined by calculating the distance traveled by the robot $\tau_{n}$.
Subsequently, for the input point cloud sequence, we use a point-based implicit representation to encode scene geometry. Starting from the first frame, the system generates samples and labels to optimize the neural implicit map, yielding a sub-map representation. From this implicit sub-map, we extract: the scene’s occupancy grid, sample points (mesh vertices) and surface normals at these points.
Next, we remove the road surface from the mesh vertices and apply Euclidean clustering to the remaining area. For each cluster, we design a feature extractor based on angular distributions between adjacent normals, deriving multiple descriptors. Following clustering of these descriptors, we obtain the geometric descriptor $\mathbf{F^{G}}\!\in\!\mathbb{R}^{a}$. Concurrently, we project the occupancy grid into a BEV, transform it to polar coordinates, and compute local responses via Log-Gabor filtering. Keypoints are selected by maximum response, with their descriptors forming the BEV descriptor $\mathbf{F^{B}}\!\in\!\mathbb{R}^{b}$. Finally, we concatenate both descriptors into the fused representation $\mathbf{F^{fuse}}\!\in\!\mathbb{R}^{a+b}$.
III-B Implicit Submap Representation With Elastic Neural Points
We will employ a memory-efficient implicit neural representation to store the geometry of the sub-map. Based on elastic neural points, this representation possesses local rotational and translational invariance. Additionally, it is unaffected by the density of the original input point cloud, has a uniform distribution, and can remove interference from certain dynamic objects.
  1. 1)
    Elastic Implicit Neural Point Cloud: The results of the sub-map composed of neural implicit points $\mathcal{M}$ are shown as follows:
    =$\mathcal{M}=\left\{\mathbf{m_{i}}=\left(\mathbf{p_{i}},\mathbf{q_{i}},\mathbf{f_{i}},t_{i}^{c},t_{i}^{u},\mu_{i}\right)\mid i=1,\dots,N\right\}$ (1)
    Let $\mathbf{p_{i}}\in R^{3}$ and the quaternion $\mathbf{q_{i}}\in R^{4}$ represent the position and orientation of each neural point $\mathbf{m_{i}}$ within the global coordinate frame, respectively. The optimizable latent feature encoding $\mathbf{f_{i}}\in R^{F}$ of each neural point serves to capture its local geometric characteristics. To efficiently track the state of every neural point, each $\mathbf{m_{i}}$ retains a creation timestamp $t_{i}^{c}$, a last update timestamp $t_{i}^{u}$, and a stability metric $\mu_{i}$.
  2. 2)
    SDF Decoder: The joint feature coordinate encoding for scene representation is employed to achieve fast convergence and reconstruct surface holes. To ensure invariance of the prediction to the local translation and rotation of neural points, the query position is transformed from the global frame to the local frame of each neural point. Specifically, let $\mathbf{d_{j}}\in\mathbb{R}^{3}$ denote the query position in the local frame of neural point $\mathbf{m_{j}}$. This local coordinate $\mathbf{d_{j}}$ is obtained by transforming the global coordinates $\mathbf{p}_{s}$ of the query position using the pose of neural point $\mathbf{m_{j}}$, as follows:
    $\mathbf{d_{j}}=\mathbf{q_{j}}(\mathbf{p_{s}}-\mathbf{p_{j}})\mathbf{q_{j}^{-1}},j=1,\dots,K$ (2)
    At the query position, $\mathbf{p_{s}}$, $\mathbf{K}$ nearby neural points are retrieved from the neural point map. For each neural point $\mathbf{m_{j}}$ in the $\mathbf{K}$-neighborhood $\mathcal{N}_{P}$, the coordinate is computed using the above Eq 2, and then the weight $\omega_{j}$ of each neural point is defined as:
    $\omega_{j}=\frac{\left\|\mathbf{d_{j}}\right\|^{-1}}{\sum_{k\in\mathcal{N}_{P}}\left\|\mathbf{d_{k}}\right\|^{-1}}=\frac{\left\|\mathbf{p_{s}}-\mathbf{p_{j}}\right\|^{-1}}{\sum_{k\in\mathcal{N}_{P}}\left\|\mathbf{p_{s}}-\mathbf{p_{k}}\right\|^{-1}}$ (3)
    Subsequently, the coordinate encoding $\mathbf{g_{i}}\in R^{C}$ or each neural point $\mathbf{m_{j}}$ is defined as $\mathbf{g_{i}}=\gamma(\mathbf{d_{j}})$, where $\gamma(\cdot)$ represents the positional encoding function. The feature encoding $\mathbf{f_{i}}\in R^{F}$ and coordinate encoding $\mathbf{g}\in R^{C}$ of the query point are derived by fusing the encodings of the nearest neighboring points along with their corresponding weights:
    $\mathbf{f}=\sum_{j\in\mathcal{N}_{P}}\omega_{j}\mathbf{f_{j}},\mathbf{g}=\sum_{j\in\mathcal{N}_{P}}\omega_{j}\mathbf{g_{j}}$ (4)
    Finally, the neural decoder $D_{\theta}$ generates the SDF prediction $s$ or the query position based on these encodings:
    $s=D_{\theta}(\mathbf{f},\mathbf{g})$ (5)
We employ a dynamic point filtering method based on the stability $\mu_{i}$ and SDF values of neural points. When updating the submap, a point is classified as a dynamic point and filtered out if its SDF value exceeds a dynamic distance threshold $\gamma_{d}$ and its stability value exceeds a stability threshold $\gamma_{\mu}$, Specifically, a sampled point $\mathbf{p_{W}}$ in the world coordinate system is considered dynamic if it satisfies the condition $\mathbf{S}(\mathbf{p_{W}})>\gamma_{d}$ and $\mathbf{H}(\mathbf{p_{W}})>\gamma_{\mu}$, where $\mathbf{S}(\mathbf{p_{W}})$ represents the SDF value indicating the distance from the point to the nearest surface, and $\mathbf{H}(\mathbf{p_{W}})$ represents the stability value quantifying the point’s consistency across multiple observations. This method effectively removes dynamic objects from the map, enhancing its accuracy and consistency.
III-C Normal Vector Descriptor
After obtaining the implicit representation of the submaps, we extract the scene’s occupancy information, the mesh vertices, and their corresponding normal vectors.
The process begins by discretizing the submap into a voxel grid with a predefined resolution $r$. At each voxel corner point within this grid, we query the Signed Distance Function (SDF) value according to the method described previously. Using these SDF values, we apply an enhanced Marching Cubes (MC) algorithm to reconstruct a triangular mesh. The MC algorithm generates vertices uniformly distributed on the reconstructed surface. The normal vector is then calculated for each vertex, defining the local surface orientation. Next, we filter the vertices to remove isolated outliers. This step refines the mesh by retaining only significant vertices along with their normals. Finally, occupancy information is derived by analyzing the SDF values at the corners of each voxel. Voxels exhibiting a sign change in their corner SDF values are classified as occupied. This occupancy classification is essential for distinguishing between occupied and free space in the environment.
We use the GPF [72] algorithm to filter out the road surface points from the input mesh vertices. This is because the road surface points contain little geometric information and provide little benefit for extracting feature descriptors for place recognition. Subsequently, inspired by the work of FEC [73], we apply a fast clustering method for mesh vertices. This method takes the mesh vertices from which the ground points have been removed as input, and the algorithm is as follows:
Algorithm 1 Fast triangular Point Cloud Clustering
1:mesh vertices point cloud $\mathbf{cloud}$, minimum cluster size $min\_size$, distance tolerance $tol$
2:Cluster labels for each point
3:Build KD-tree $\mathbf{T}$ from $\mathbf{cloud}$
4:$n\leftarrow|\mathbf{cloud}|$
5:if $n<min\_size$ then
6:     throw ”Point cloud too small”
7:end if
8:Initialize $labels[0..n-1]\leftarrow 0$
9:$current\_label\leftarrow 1$
10:for each point $p_{i}\in cloud$ do
11:     if $labels[i]\neq 0$ then
12:         continue
13:     end if
14:     $neighbors\leftarrow\mathbf{T}.\texttt{radius\_search}(p_{i},tol)$
15:     if $neighbors$ is empty then
16:         continue
17:     end if
18:     $min\_tag\leftarrow\infty$
19:     for each $j\in neighbors$ do
20:         if $labels[j]>0$ then
21:              $min\_tag\leftarrow\min(min\_tag,labels[j])$
22:         end if
23:     end for
24:     if $min\_tag=\infty$ then
25:         $min\_tag\leftarrow current\_label$
26:         $current\_label\leftarrow current\_label+1$
27:     end if
28:     for each $j\in neighbors$ do
29:         if $labels[j]>min\_tag$ then
30:              Relabel all points with $labels[k]=labels[j]$ to $min\_tag$
31:         end if
32:         $labels[j]\leftarrow min\_tag$
33:     end for
34:end for
Following the derivation of clusters $\mathbf{\mathcal{C}=\left\{C_{1},...,C_{M}\right\}}$, points within each cluster are normalized. This normalization reduces sensitivity to noise and improves the robustness of subsequent geometric feature extraction. Specifically, all points in a cluster are isotropically scaled to be contained within a unit sphere of radius 1.
Next, the spherical domain is partitioned into 72 sub-regions based on longitude and latitude. The sphere is first divided along the z-axis into northern and southern hemispheres. By latitude, each hemisphere is further divided into two bands, resulting in 4 latitude bands globally. Longitudinal division then subdivides each latitude band into 18 sectors, formed by meridians spaced 20 degrees apart. Fig. 3 illustrates this spatial subdivision.
Figure 3: The figure illustrates the extraction of geometric descriptors for 3D segments. The input is a 3D segment with surface normals; color coding indicates normal deviations. First, these normals are mapped to discrete bins on a unit sphere. Then, the average angular deviation of the normals is calculated within each bin. Finally, these per-bin statistics are assembled into the geometric descriptors.
Following the spherical partitioning, the mean normal vector is computed for each sub-region within a cluster. The angular deviation between all unique pairs of these regional mean normal vectors within the same cluster is then calculated, generating $\binom{n}{2}$ distinct angular measurements per cluster (where $n$ is the number of sub-regions, e.g., $n=72$). A histogram of these angular deviations is constructed per cluster using 10-degree bin intervals (resulting in $d=18$ dimensions), yielding a $d$-dimensional geometric descriptor vector $\mathbf{F_{C}}\in\mathbb{R}^{d}$ for each cluster $C$.
Subsequently, for each submap, the geometric descriptors $\mathbf{F_{C}}$ from all its clusters are collected and normalized. The VLAD algorithm [74] is then applied to this normalized set of descriptors to generate a single submap-level geometric descriptor. To accommodate submaps with varying cluster counts ($N_{c}$), we dynamically adjust the size of the VLAD codebook (i.e., the number of cluster centers, $K$) during encoding, setting $K$ proportionally to $N_{c}$. Submaps with a small $N_{c}$ use a correspondingly smaller $K$, while those with a large $N_{c}$ use a larger $K$. The resulting VLAD vector is then zero-padded to achieve a fixed dimensionality $a$. Finally, this yields the standardized geometric descriptor $\mathbf{F^{G}}\in\mathbb{R}^{a}$ for the submap. The entire process is illustrated in Fig. 3.
III-D BEV descriptor
  1. 1)
    Log-Gabor filters: We project the acquired occupancy grid information onto the z-plane to obtain a BEV. The size of each pixel in the BEV corresponds to the physical size of the occupancy grids. Subsequently, we normalize the pixel values of the input image to the range [0, 255] to ensure numerical stability in subsequent processing. Then, we convert the input image’s coordinates (u,v) into polar coordinate representation $\left(\rho,\theta\right)$.Subsequently, we construct a Log-Gabor filter [75], defined as follows:
    $L\left(f,\omega,s,o\right)=exp(-\frac{(log(f/f_{s}))^{2}}{2(log(\sigma_{f}/f_{s}))^{2}}-\frac{(\omega-\omega_{0})^{2}}{2\sigma_{\omega}^{2}})$ (6)
    Among them, $f_{s}$ and $\omega_{o}$ are the center frequency and direction of the filter, respectively. $\sigma_{f}$ and $\sigma_{\omega}$ are the width parameters of the filter. Calculate each pixel point’s filter responses at different scales and directions using the Log-Gabor filter. First, we perform the convolution operation, which involves convolving the filter with the input image:
    $A\left(\rho,\theta,s,o\right)=\left\|B\left(\rho,\theta\right)\ast L(\rho,\theta,s,o)\right\|^{2}$ (7)
    Subsequently, for each direction $o$, the responses of all scales $s$ are aggregated:
    $A(\rho,\theta,o)=\sum_{s}A(\rho,\theta,s,o)$ (8)
  2. 2)
    Maximum Index Map (MIM): MIM is an orientation map that records the direction of the maximum response of each pixel in all directions [76]. That is, for each pixel $(\rho,\theta)$, find the direction $o$ with the maximum response:
    $MIM(\rho,\theta)=arg\max_{o}A(\rho,\theta,0)$ (9)
    The size of the MIM is the same as that of the input image, and each pixel stores a direction index. Subsequently, the FAST corner detection algorithm is utilized on the MIM to extract key points. Subsequently, for each key point, we calculate the dominant direction around it. A MIM region of size $J\times J$ is taken around the key point, and an orientation histogram $\mathbf{h(o)}$ is constructed for this region to count the pixel values in each direction. Then, find the peak direction $o_{m}$ of the histogram:
    $o_{m}=arg\max_{o}\mathbf{h(o)}$ (10)
    The main direction $\beta$ is:
    $\beta=\frac{\pi o_{m}}{N_{o}}$ (11)
    Align the direction of the MIM region concerning the main direction to achieve rotational invariance, and rotate the MIM region by an angle of $\beta$.
    $patch_{\beta}(\rho,\theta)=mod(patch(\rho,\theta+\beta)-o_{m},N_{o})$ (12)
    The adjusted MIM region is partitioned into $l\times l$ sub-grids. A directional histogram is constructed for each sub-grid to count the pixel values in each direction. The histograms of all sub-grids are concatenated to form a feature vector. Finally, all the feature vectors are processed using the VLAD method to obtain the final BEV descriptor $\mathbf{F^{B}}\!\in\!\mathbb{R}^{b}$.
IV EXPERIMENTAL RESULTS AND ANALYSES
In this section, Section IV-A describes the dataset selected by us and some experimental settings, Section IV-B introduces the evaluation metrics of the experiment, Section IV-C presents the experimental results. Section IV-D and Section IV-E show the ablation experiment and running time situation respectively.
IV-A Dataset and Experimental Settings
Below, we provide a detailed description of the dataset used in this study. Our dataset includes various types of LiDAR and different application scenarios. The configuration of the LiDAR sensors used in this dataset is shown in Fig. 4.
Figure 4: Datasets for evaluation.
  1. 1)
    KITTI Dataset [77]: This dataset was collected at 10 Hz in an urban environment using a mechanical LiDAR (Velodyne HDL-64E) mounted on top of a vehicle. We selected six sequences (00, 02, 05, 06, 07, 08) containing loop closure points for our experiments.
  2. 2)
    KITTI-360 Dataset [78]: This dataset was also collected at 10 Hz in an urban scenario using a LiDAR (Velodyne HDL-64E). We selected four sequences (00, 04, 05, 06) containing loop closure points for the experiment.
  3. 3)
    NCLT Dataset [79]: This dataset was collected using a 32-line LiDAR (Velodyne HDL-32E) mounted on a Segway robot. It was primarily collected on the University of Michigan campus, capturing variations across the four seasons, differing lighting conditions, and dynamic objects. Based on loop closure availability, we selected four sequences for the experiment: 2012-05-26 (NCLT01), 2012-08-20 (NCLT02), 2012-09-28 (NCLT03), and 2013-04-05 (NCLT04). Additionally, sequences 2012-03-17, 2012-02-04, and 2012-08-20 were used for cross-temporal location recognition experiments.
  4. 4)
    MulRan Dataset [80]: This dataset contains data collected from multiple scenarios in South Korea using an Ouster OS1-64 LiDAR. We selected sequences 01 and 02 from the DCC, KAIST, and Riverside scenes for the experiment; these were labeled MulRan01 to MulRan06 respectively (e.g., DCC01=MulRan01, DCC02=MulRan02, etc.). Additionally, sequences Sejiong01 to Sejiong03 were used to conduct cross-temporal location recognition experiments.
Our feature extraction method is manually applied to the sub-maps derived from the accumulated point clouds. To benchmark our approach, we selected five handcrafted feature extraction methods for comparison on the evaluated datasets. These methods are: Scan Context++ [56], M2DP [54], NDT [81], BoW3D [82], and STD [18]. Among them, M2DP, NDT, BoW3D, and STD can be applied directly to the accumulated sub-maps. To ensure a fair comparison, we also applied the Scan Context++ method to the sub-maps. For this purpose, we re-projected the accumulated sub-map points onto the middle frame of the sub-map to simulate a dense ”scan” captured from that pose. For BoW3D, as its default configuration is tailored to the KITTI dataset, we only conducted experiments on KITTI and KITTI-360. Experiments involving our method and five handcrafted feature-based baseline approaches were executed on a laptop equipped with an Intel Core i7-10875H CPU @ 5.10GHz and 16GB RAM. For constructing the neural implicit representations, we used an RTX 2060 GPU for acceleration. The parameter values used in our method are listed in Table I.
TABLE I: Parameters of our methold
Parameter Value Description
$\tau_{n}$ 25m movement distance to form a submap
$K$ 8 neighborhood neural point number
[…]
$r$ 0.2 voxel resolution
$min\_size$ 200 minimum number of clusters
$l\times l$ $6\times 6$ number of subgrids
$J$ 96 local patch size
IV-B Performance Evaluation Metrics
In this article, the conditions for establishing a loop closure are defined as follows: For a query sub-map, the average position of all its frames is computed and regarded as the position of that sub-map. If the distance between this sub-map’s position and that of a previous sub-map is less than or equal to 20 meters, and the difference in their indices is greater than 50, then these two sub-maps are considered positive (loop closure) pairs. The distance threshold of 20 meters was chosen based on considerations of the urban environment layout, typical LiDAR sensor range, and overall scale. Each method outputs a candidate frame along with its similarity score and compares the score against its predefined decision threshold. If the score exceeds the threshold, the match is classified as positive; otherwise, it is classified as negative. For a match classified as positive, the actual geometric distance between the query frame and the candidate frame is computed. If this distance is less than 20 meters, it is classified as a true positive (TP); otherwise, it is classified as a false positive (FP). For a match classified as negative: If no ground truth loop closure exists for the query frame (according to the criteria above), it is classified as a true negative (TN). However, if a ground truth loop closure does exist for the query frame, it is classified as a false negative (FN).
The following is how the evaluation metrics we adopt are calculated:
  1. 1)
    Precision–Recall Curve: In the field of place recognition, precision is defined as the ratio of true positive results to the total number of identified matches. Recall, in contrast, represents the ratio of true positive results to the total number of actual positive instances. These concepts can be formally expressed as:
    $Precision=\frac{TP}{TP+FP}$ (13)
    $Recall=\frac{TP}{TP+FN}$ (14)
    Here, TP denotes the number of true positive outcomes, FP represents the number of false positive results, and FN indicates the number of false negative instances. The precision-recall curve is constructed by adjusting decision thresholds. Each point on this curve corresponds to a specific threshold setting, illustrating the precision-recall trade-off and visually demonstrating how these metrics co-vary with threshold adjustments.
  2. 2)
    AUC: The Area Under the Curve (AUC) quantifies the area beneath the Precision-Recall (PR) curve. It serves as a widely adopted metric for evaluating model performance across various decision thresholds. As a single numerical value, AUC ranges from 0 to 1. A higher AUC score indicates superior performance, reflecting the model’s enhanced capacity to accurately classify positive instances across different threshold settings.
  3. 3)
    Max F1 Score: The F1 score, calculated using Eq 15, represents the harmonic mean of precision and recall. This score provides a balanced measure that combines both metrics. The Maximum F1 score serves as a comprehensive performance indicator, often used to identify the optimal decision threshold. By balancing the often-competing demands of precision and recall, it effectively reconciles these priorities. The peak F1 score value indicates the algorithm’s overall best performance, demonstrating an optimal equilibrium between correctly identifying true positives and minimizing false positives.
    $F_{1}=2\times\frac{Precision\times Recall}{Precision+Recall}$ (15)
Figure 5: Evaluation on twenty short-term sequences reveals that each subfigure illustrates the precision-recall performance of different methods. Our approach consistently outperforms all others across these sequences, demonstrating its robustness and adaptability.
TABLE II: AUC AND MAX F1 SCORES ON TESTING SEQUENCES
Method KITTI00 KITTI02 KITTI05 KITTI06 KITTI07 KITTI08 KITTI360-00 KITTI360-04 KITTI360-05 KITTI360-06
M2DP 0.64/0.65 0.31/0.45 0.55/0.47 0.52/0.58 0.28/0.43 0.43/0.58 0.48/0.56 0.08/0.23 0.09/0.28 0.18/0.31
NDT 0.69/0.70 0.52/0.59 0.61/0.67 0.6/0.68 0.38/0.52 0.52/0.6 0.51/0.57 0.26/0.44 0.44/0.56 0.39/0.48
SC++ 0.72/0.75 0.63/0.67 0.55/0.65 0.69/0.73 0.47/0.60 0.27/0.44 0.61/0.64 0.58/0.47 0.35/0.45 0.42/0.48
BoW3D 0.78/0.80 0.74/0.74 0.78/0.76 0.55/0.63 0.66/0.70 0.63/0.64 0.67/0.68 0.79/0.8 0.71/0.7 0.57/0.62
STD 0.51/0.58 0.85/0.8 0.78/0.76 0.71/0.77 0.57/0.62 0.73/0.74 0.71/0.71 0.6/0.64 0.68/0.71 0.63/0.64
Ours 0.93/0.86 0.95/0.88 0.93/0.91 0.91/0.89 0.91/0.82 0.83/0.78 0.89/0.86 0.87/0.85 0.87/0.86 0.76/0.77
Method NCLT01 NCLT02 NCLT03 NCLT04 MulRan01 MulRan02 MulRan03 MulRan04 MulRan05 MulRan06
M2DP 0.37/0.44 0.58/0.62 0.47/0.57 0.49/0.53 0.10/0.30 0.07/0.21 0.10/0.23 0.02/0.10 0.02/0.13 0.09/0.23
NDT 0.75/0.69 0.18/0.26 0.57/0.61 0.51/0.61 0.41/0.55 0.21/0.38 0.17/0.34 0.13/0.30 0.18/0.35 0.15/0.30
SC++ 0.74/0.74 0.67/0.64 0.63/0.69 0.44/0.52 0.41/0.55 0.36/0.48 0.34/0.48 0.24/0.42 0.29/0.44 0.17/0.32
STD 0.61/0.65 0.51/0.62 0.74/0.75 0.64/0.70 0.55/0.62 0.43/0.55 0.36/0.50 0.34/0.48 0.49/0.63 0.32/0.44
Ours 0.92/0.85 0.77/0.78 0.83/0.81 0.82/0.78 0.70/0.73 0.70/0.69 0.50/0.58 0.40/0.56 0.59/0.67 0.47/0.55
Note: On the left is the AUC and on the right is the max F1 score. The best-performing method is bolded.
Figure 6: Loop retrieval result using each method’s max F1 score threshold on KITTI02.
Figure 7: Loop retrieval result using each method’s max F1 score threshold on MulRan04.
IV-C Experiment Performance Comparisons
To comprehensively test our method, we evaluated its performance in two major scenarios: short-term and long-term relocalization. Here, ”short-term” refers to real-time detection of loop closure points, while ”long-term” denotes the ability to recognize locations when revisiting the same area after a long period, specifically more than one day.
  1. 1)
    short-term: The PR curves for our proposed method and existing methods are shown in Fig. 5. We compared a total of 20 trajectories. Our method demonstrates superior performance compared to the existing methods. In the KITTI00 sequence, since there are only small rotational and translational changes when passing through the same location, most methods perform well. However, in scenarios with significant rotational and translational changes at the same location – such as KITTI05, KITTI08, KITTI360_04, and KITTI360_05 – existing methods tend to generate inaccuracies, whereas our method retains robust performance. When dealing with numerous similar/repetitive scenes or revisiting the same location multiple times (e.g., the six sequences in the MulRan dataset), all methods show some susceptibility to these challenging conditions. Nevertheless, our method still maintains good performance. Table II presents the AUC values and peak F1 scores. Our method surpasses the compared approaches. To provide a detailed analysis of each method, two sets of sequences for testing are presented in the figure. All methods display their optimal performance (indicated by the maximum F1 score). Compared to other methods, ours achieves higher True Positives (TP) and lower False Positives (FP) and False Negatives (FN). To further highlight the distinction, we selected KITTI02 and MulRan04 tracks and visualized their place recognition results (including TP, FP, FN, and TN) based on each method’s optimal F1 score threshold. As illustrated in Figs. 6 and 7, it is evident that our approach generates fewer false matches and better detects loopback points, thereby delivering superior performance.
  2. 2)
    long-term: A robust place recognition system should perform reliably despite long-term environmental variations. To evaluate long-term place recognition performance, multiple intermittent datasets covering the same area were selected as the map sequence and query sequence, respectively. The results of our recall-precision experiments are shown in Fig. 8. AUC and F1 scores are presented in Table III. Our method demonstrates superior performance across all metrics. Although performance declines compared to short-term experiments on the same dataset, this results from the expanded retrieval scope and environmental changes inherent to long-term tasks, which increase place recognition complexity.
As demonstrated, our method outperforms comparative approaches. The key factors contributing to this superior performance are threefold. First, our implicit representation ensures enhanced uniformity in the acquired BEV, which directly translates to more consistent spatial descriptors. Second, our geometric descriptors strategically disregard ground-level information (prone to redundancy) while focusing exclusively on regions containing rich geometric features such as edges and corners—a distinct methodological departure. Finally, the implicit representation inherently suppresses dynamic elements, enabling effective noise interference elimination.
TABLE III: AUC AND MAX F1 SCORES ON TESTING SEQUENCES.
Method NCLT-L1 NCLT-L2 MulRan-L1 MulRan-L2
M2DP 0.23/0.35 0.25/0.39 0.09/0.22 0.08/0.25
NDT 0.30/0.39 0.31/0.45 0.22/0.38 0.10/0.23
SC++ 0.49/0.57 0.39/0.49 0.29/0.45 0.18/0.34
STD 0.51/0.54 0.50/0.57 0.33/0.45 0.27/0.41
Ours 0.75/0.77 0.70/0.73 0.38/0.51 0.35/0.48
Note: On the left is the AUC and on the right is the max F1 score. The best-performing method is bolded. ”NCLT-L1” and ”NCLT-L2” represent ”2012-03-17” to ”2012-02-04” and ”2012-08-20” to ”2012-02-04”, respectively. MulRan-L1”, and ”MulRan-L2” represent ”Sejiong02” to ”Sejiong01 ”, and” Sejiong03” to ”Sejiong01 ”, respectively.
Figure 8: Precision–recall curve NCLT, MulRan on long-term.
TABLE IV: AUC AND MAX F1 SCORES ON ABLATION STUDY.
Implicit Representation BEV Image Normal Vector KITTI KITTI-360 NCLT Mulran
0.91/0.86 0.85/0.84 0.8/0.79 0.51/0.60
0.59/0.57 0.50/0.48 0.47/0.42 0.32/0.27
0.76/0.79 0.75/0.77 0.73/0.71 0.42/0.40
0.81/0.79 0.72/0.73 0.76/0.78 0.48/0.50
Note: On the left is the AUC and on the right is the max F1 score. The best-performing method is bolded. The individual datasets are presented as their average levels.
IV-D Ablation Study
To analyze the contributions of individual method components to place recognition performance, we designed ablation experiments. Specifically, we examined whether including the implicit neural representation, as well as the BEV and normal vector modules, affected 3D place recognition accuracy.
To validate the contribution of the point-based implicit representation to our place recognition framework, we conducted two experimental configurations on our dataset. The first configuration retained all default parameters, while the second removed the implicit representation module. In the absence of an implicit submap, we adopted a voxelization-based statistical method to generate occupancy grids for standard submaps. Specifically, we calculated point density within each voxel and applied a density threshold matching that of implicit submaps to ensure resolution consistency. For normal vector estimation, we employed a principal component analysis (PCA)-based approach: PCA was computed on each point’s local neighborhood, with the eigenvector corresponding to the smallest eigenvalue designated as the surface normal. Ablation tests were performed across four datasets covering short-term and long-term scenarios. Quantitative results (AUC and Maximum F1-score) for all datasets are summarized in Table IV.
Building on the implicit submap representation, we conducted two additional experiments: one utilizing exclusively BEV features and the other using solely normal vector information features. Detailed quantitative comparisons are provided in Table V. As evidenced in the table, removing the implicit submap representation causes marked performance degradation. This decline stems primarily from two sources: 1) errors in surface normal estimation, and 2) inaccuracies in occupancy grid generation, both of which introduce noise that compromises 3D place recognition robustness. It can also be found that only the macro-level information from the BEV perspective and the local micro-scale information on the main 3D segments cooperate to produce a robust descriptor.
TABLE V: TIME CONSUMPTION (MS) OF DIFFERENT METHOD.
Method KITTI KITTI360 NCLT MulRan
M2DP 73.2 74.1 80.2 76.2
NDT 465.2 451.2 387.2 392.1
SC++ 62.1 59.4 57.2 63.1
STD 38.2 33.2 36.9 40.1
Ours 60.2 71.3 63.4 62.8
Ours(no implicit) 31.6 37.2 33.9 35.2
Ours(only implicit) 28.6 34.2 29.5 27.6
IV-E Computational Cost
We conducted comparative experiments on the runtime of our method to demonstrate its efficiency in 3D place recognition. We statistically analyzed the average time required to generate descriptors for each submap across all sequences, including calculating the average time for the three most critical components, and compared our method with other approaches. The running time of our method was calculated by measuring the time for descriptor generation and retrieval. The results are presented in Table V. From the table, we can see that our method is not worse than most methods in running time, even if we use implicit expressions. Our implicit representation is lightweight and efficient. We add the Voxel Hashing technique to the process of obtaining nerve points to index nerve points efficiently and reduce some unnecessary optimization iterations. Meanwhile, our method for segmenting major 3D segments is also fast and efficient. Finally, real-time performance is achieved.
V CONCLUSION
In this paper, we propose a novel 3D place recognition method inspired by implicit representations. Leveraging submap-based implicit representations, the method begins by acquiring only occupancy grids, uniform triangular points, and corresponding globally consistent normal vectors. This information is then collaboratively utilized to generate BEV images and information on clustered 3D segments, depending on exact meaning. We employ Log-Gabor filters to derive descriptors for the BEV images and compute descriptors for the 3D segments based on the angular differences between their normal vectors. By concatenating these two descriptors, we construct a cross-modal geometric hierarchy fusion feature. To the best of our knowledge, this is the first framework that utilizes implicit representations to fuse cross-modal geometric features for 3D place recognition, significantly advancing the exploitation of implicit representations’ potential in 3D place recognition. In future work, we will further explore the possibilities and roles of implicit representations in place recognition, including the development of new representation paradigms specifically tailored for this task.